OpenStudy (anonymous):

Simplify the expression: (92wy)^3 x 3wy

OpenStudy (anonymous):

\(\large (92wy)^3 \cdot 3 \cdot w \cdot y \) = \(\large (92wy) \cdot(92wy) \cdot(92wy) \cdot 3 \cdot w \cdot y \) = \(\large 92 \cdot w \cdot y \cdot 92 \cdot w \cdot y \cdot 92 \cdot w \cdot y \cdot 3 \cdot w \cdot y \) = \(\large 92^3 \cdot 3 \cdot w^4 \cdot y^4\)